OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive?

OpenStudy (anonymous):

if Sally received 25% of 125,000and lost by 62,500 then Maya must have: x = 0.25(y) + 62,500 because Maya has 62,500 more than sally and sally has 25%. Plugging in the 125,000 we get: x= 0.25(125,00) + 62,500 x= 31,250 + 62,500 x= 93,750 votes assuming my arithmatic was correct